# Vib2Mol: from vibrational spectra to molecular structures—a versatile deep learning model

### v2

@@ Abstract
There will be a paradigm shift in chemical and biological research, to be enabled by autonomous, closed-loop, real-time self-directed decision-making experimentation. Spectrum-to-structure correlation, which is to elucidate molecular structures with spectral information, is the core step in understanding the experimental results and to close the loop. However, current approaches usually divide the task into either database–dependent retrieval and database–independent generation and neglect the inherent complementarity between them. In this study, we proposed Vib2Mol, a general deep learning model designed to flexibly handle diverse spectrum-to-structure tasks according to the available prior knowledge by bridging the retrieval and generation. It achieves state-of-the-art performance, even for the most demanding Raman spectra, over previous models in predicting reaction products and sequencing peptides as well as analyzing experimental spectra and integrating multi-modal spectral data. Vib2Mol enables vibrational spectroscopy a real-time guide for autonomous scientific discovery workflows.111codes are available at https://github.com/X1nyuLu/vib2mol
@@ 1 Introduction
In this study, we propose a DL-based vibrational spectrum-to-molecular structure model (Vib2Mol) to flexibly address a variety of spectral annotation tasks according to the available prior knowledge. We focus particularly on Raman spectroscopy, a widely used non-invasive, in-situ method. Raman spectroscopy has long faced challenges in accurately simulating the Raman spectra of target species, which requires very demanding calculation of high-order energy and force derivative. It becomes even challenging to achieve satisfactory alignment between theoretical and experimental spectra and interpret the experimental spectra. If we can demonstrate the feasibility of Vib2Mol for such intricate form of spectra, if may be easily extended to other spectroscopic techniques. In specific, Vib2Mol adopts an encoder-decoder transformer architecture, and is trained with the strategy of multi-task learning and integrates a wide variety of spectrum-to-structure tasks into one versatile model. Our model achieves state-of-the-art performance on 9 out of 10 test sets when compared to mainstream DL-based methods. It further enhances the accuracy in interpreting spectra of reaction products and peptide sequencing as more knowledge of target molecules is introduced. Additionally, it demonstrates robustness in analyzing experimental spectra and potential for integrating multi-modal spectra from multiple techniques. This advancement demonstrates significant potential for in-situ intelligent analysis of dynamic chemical transformations and biological processes.
@@ 2.1 Multi-task learning framework: correlating vibrational spectrum and molecular structure
The workflow of Vib2Mol during pre-training, including alignment and generation modules is illustrated in Figure 1. The alignment module (Figure 1A) aims to bring the spectral and structural features of the same molecule as close as possible while separating the features of different molecules simultaneously. Spectra and molecular structures are represented as patch tokens and SMILES tokens, and then encoded into spectral and molecular embeddings by encoders, respectively. These two embeddings are effectively aligned through contrastive learning (CL), enabling cross-modal spectrum-structure retrieval. Since spectrum-spectrum retrieval is implicitly included in the spectrum-structure contrastive learning, there is no need to explicitly design a specific loss function dedicated to spectrum-spectrum retrieval only.
@@
Figure 1B depicts the workflow of conditional generation and de novo generation of molecular structures. Conditional generation, i.e., predicting the occluded molecular structure on the basis of the spectrum, draws on masked language modeling (MLM). Briefly, SMILES tokens, representing the molecular structure, are randomly masked by 45% and then processed by molecular encoders to generate molecular features. Note that molecular encoders share parameters with the alignment module in Figure 1A. Then molecular decoders fused information from both masked molecular embeddings and spectral features, and predicted the to-be-determined tokens using cross-attention. Differently, De novo generation draws on language modeling (LM). Briefly, SMILES tokens are sequentially masked from left to right and directly input into the molecular decoders sharing parameters with MLM. Guided by spectral features and previously generated SMILES sequences, the decoders can predict the next SMILES token from left to right until the entire sequence is complete.
@@
Figures 1C to 1F illustrate the workflow of Vib2Mol during application and inference. (1) For spectrum-spectrum retrieval (Figure 1C), instead of directly comparing spectral similarity by metrics such as Pearson correlation coefficient, the to-be-determined spectrum is encoded into an embedding vector, and the cosine similarity is calculated between this vector and the known spectral embedding vector in the database. (2) For spectrum-structure retrieval task (Figure 1D), Vib2Mol uses CL to minimize the distance between the spectral and structural features of the same molecule, referring to the mainstream spectrum-structure contrastive learning models[3, 5, 15]. spectrum-structure retrieval is based on spectrum-structure similarity, i.e., the cosine value between the spectral embedding of the to-be-determined spectrum and the molecular embeddings of known molecules in the database. (3) For conditional generation task (Figure 1E), Vib2Mol adopts the encoder-decoder architecture. Both the spectrum and partially masked molecular structure are encoded and then fused through the molecular decoder to generate the SMILES of the masked part. (4) For de novo generation (Figure 1F), Vib2Mol directly employs molecular decoders to predict SMILES one by one on the basis of the encoded spectral features until a complete molecular structure is generated. When generating the next token, stochastic perturbation beam search (BS, see Methods for details) is used to ensure the diversity of the results, which is also confirmed to improve the generation performance by ablation experiments (Table S1).
It is worth noting that Vib2Mol adopts staged pre-training (SPT), the advantage of which is confirmed by ablation experiments in Table S1. In the first stage, the alignment module (including the spectral and molecular encoders) is trained by CL loss. After that, the parameters of the trained spectral and molecular encoders are frozen. In the second stage, the generation module (molecular decoder) is trained by MLM and LM losses. The molecular encoder of Vib2Mol shares parameters for CL and MLM, and the molecular decoder shares parameters for MLM and LM. Meanwhile, the spectral features extracted by the spectral encoder are reused all the time. It is the parameter sharing and feature reusing strategies that enables Vib2Mol to address above spectrum-to-structure tasks by simply changing the combination of encoders and decoders, without the need for additional fine-tuning or training.
@@
Figure 1: (A-B) The architecture of Vib2Mol for pretraining and (C-F) the workflow for different spectrum-to-structure tasks. (A) The alignment module: spectra and molecular structures are represented as patch tokens and SMILES tokens, respectively. After processed by their encoders, spectral and molecular information are aligned by CL. (B) The generation module: for conditional generation, molecules are randomly masked 45% and encoded by the same molecular encoder used for spectrum-structure alignment. The molecular decoder fuses spectral information with molecular features and predicts masked tokens. MLM loss is used to compare the distance between predicted and original SMILES. For de novo generation, molecule is sequentially masked and directed input into the same molecular decoder as conditional generation without the prior encoding. Then, the decoder predicts the next token based on previous information and spectral features. By flexibly combining encoders and decoders, Vib2Mol can address four spectrum-to-structure tasks at the same time: (C) spectrum-spectrum retrieval, where only the spectral encoder is used to calculate the similarity between spectral pairs; (D) spectrum-structure retrieval, where spectra and molecules are encoded by their respective encoders to determine spectrum-structure similarity; (E) conditional generation, and (F) de novo generation, both following workflows during the stage of pretraining.
@@ 2.2 State-of-the-art performance of Vib2Mol
Figure 2A shows the performance of Vib2Mol on the vibrational spectrum-to-structure benchmark (ViBench, see Methods for details), in comparison with three mainstream DL-based models: vanilla-CL, vanilla-MLM, and vanilla-LM (see methods for more details about baseline models). Vib2Mol is not only capable of addressing all tasks simultaneously but also achieves the best performance on 9 of 10 test sets (detailed metrics are recorded in Table S1).
For spectrum-spectrum retrieval, the Recall@1 of Vib2Mol (76.46%) in the VB-geometry is comparable to that of vanilla-CL (76.90%), both of which are significantly better than traditional methods (Recall@1 $\leq$ 34.71%). Such huge difference is a result of different matching strategies used for these three methods. Instead of directly using the cosine value or Pearson correlation coefficient to evaluate the similarity of spectral pairs, employing deep learning to project spectrum into the latent space is important. For spectrum-structure retrieval, on the VB-mols test sets, Vib2Mol can further increase the Recall@1 of vanilla-CL from 78.07% to 78.92%.
For the generation tasks, Vib2Mol can not only address them simultaneously but also achieve better performance. Taking the VB-mols as an example, the molecular accuracy of Vib2Mol, a metric for conditional generation (see more in Methods), reached 92.54%, better than vanilla-MLM (90.67%), and the Recall@1 of de novo generation was 56.16%, better than that of vanilla-LM (55.15%).
It is worth noting that the performance of all models on VB-qm9 is generally better than that on VB-zinc15. This is due to the much less molecular diversity of VB-qm9 than that of VB-zinc15, resulting in a smaller search space in retrieval and generation. Overall, the Vib2Mol model outperforms dedicated models designed for single tasks, indicating a synergistic effect between multiple spectrum-to-structure tasks.
2.3 Synergistic interaction among modules of Vib2Mol
The synergistic interaction was explored by recursively combining three single-task models, including spectrum-structure retrieval model trained with CL only, conditional generation model trained with MLM only, and de novo generation model trained with LM only. As shown in Figure 2B and Table S2, on the test set of VB-mols, the best model for spectrum-structure retrieval is CL-only. However, as more tasks are introduced (CL→CL+MLM→CL+MLM+LM), the retrieval performance gradually decreases (Recall@1 drops from 78.92% to 76.26% and 75.35%). In contrast, the performance in de novo generation continuously improves (Figure 2D) with the Recall@1 increasing from 54.99% (LM-only) to 57.84% (CL+LM) and then 58.68% (CL+MLM+LM). Meanwhile, for conditional generation, the accuracy increases from 93.59% to 94.03% with the introduction of CL, then decreases to 92.85% with the further introduction of LM (Figure 2C).
To further investigate the mechanism of the interaction among three tasks, taking CL+MLM+LM as an example, the relative contribution of each loss to updating parameters was quantified (Figure S1, see Methods for the detailed calculation). With the increase of epochs, the contribution of CL loss significantly decreases and approximately equals to 0% when epoch is larger than 200. In contrast, the contribution of MLM loss remains around 20-40%, and that of LM loss always dominates while gradually increasing.
From the perspective of difficulty of task, generation requires searching for molecular structures that meet multiple constraints in an open chemical space[16], while the smaller search space and more clear problem definition both offer a lower complexity of retrieval[17]. As a result, generation losses (MLM and LM) are more likely to dominate the optimization[18, 19, 20], resulting in weaker retrieval performance. At this point, CL loss with lower contribution acts similarly to auxiliary regularization and leads to stronger generation performance. Interestingly, the conditional generation metric of CL+MLM falls between MLM-only and CL+MLM+LM due to the moderate contribution of MLM. This phenomenon further confirms the order of these tasks: spectrum-structure retrieval (normal),conditional generation (harder), and de novo generation (hardest). In summary, the simpler task has lower loss contribution, promoting the performance of harder task with higher loss contribution.
Figure 2: (A) Performance of single-task models and Vib2Mol (red region) on 10 test sets. Blue, purple, and green regions represent retrieval, conditional generation, or de novo generation tasks, respectively. Ablation experiments of multi-task learning for (B) retrieval, (C) conditional generation, and (D) de novo generation.
2.4 Visualizing molecular and spectral embeddings to elucidate compound class clusters
In order to verify whether the superior performance of Vib2Mol stems from its unbiased understanding of the semantic information of molecules and spectra, the following two experiments were carried out.
From the macroscopic perspective, we randomly selected one million molecules from VB-qm9 and calculated their structural and spectral similarities. As shown in Figure 3A, on the one hand, a significant positive correlation trend (Pearson correlation coefficient is 0.65) is between the molecular similarity and spectral similarity, both of which are extracted by Vib2Mol. Such correlation indicates that the model can well align the structural and spectral information of molecules. On the other hand, there is an obvious variance between the two similarities during the range of 0-0.5, indicating that there are certain differences between them. We thereby sampled spectrum-molecule pairs containing common functional groups (ether, alcohol, amine, nitrile, amide, ketone, alkene, haloalkane, alkyne, imide). Subsequently, t-SNE was employed to visualize the distribution of structural and spectral embeddings of molecules with different functional groups. Molecules with the same functional group can be clustered to several groups either by structural embeddings (Figure 3B) or spectral embeddings (Figure 3C). In addition, the overall distribution and clustering of structural or spectral embeddings are similar, confirming the reliable performance of Vib2Mol in aligning molecular structures and spectra.
From the microscopic perspective, although similar molecular embeddings are close to each other in distance, molecules with the same functional group often gather into several clusters, due to the fact that molecular structures consist of both functional groups and backbones. For instance, although molecules 1 and 2 are classified as alcohol, the difference in backbones, in which a carbon ring is for molecule 1 and a heterocycle containing O and N is for molecule 2 (Figure 3D), brings very low molecular similarity (0.02, Figure 3E) and spectral similarity (0.23, Figure 3F) between these two molecules, respectively. In contrast, although molecule 2 and molecule 3 (nitrile) are grouped to different clusters, the much similar backbone (both are heterocycles containing O and N) offers both higher molecular similarity (0.32) and spectral similarity (0.30) between them than those between molecules 1 and 2.
In summary, at the macro level, Vib2Mol can well understand the structural and spectral information of molecules and effectively extract the relationship between them. At the micro level, a tight connection is between molecular similarity and spectral similarity, but with a certain gap. This phenomenon is not surprising at all, since Raman spectrum cannot completely reveal all the structural information, but only reflect part of the characteristics of a molecule from the aspect of vibration, resulting in a deviation between the calculated similarity and actual molecular characteristics.
Figure 3: (A) Joint distribution of molecular similarity and spectral similarity. Visualization of (B) molecular and (C) spectral embeddings using t-SNE. (D) Molecular structure and related Raman spectra of selected molecules. Similarity matrices of molecular (E) and spectral (F) embeddings.

### v3

@@ Abstract
There will be a paradigm shift in chemical and biological research, to be enabled by autonomous, closed-loop, real-time self-directed decision-making experimentation. Spectrum-to-structure correlation, which is to elucidate molecular structures with spectral information, is the core step in understanding the experimental results and to close the loop. However, current approaches usually divide the task into either database-dependent retrieval and database-independent generation and neglect the inherent complementarity between them. In this study, we proposed Vib2Mol, a versatile deep learning model designed to flexibly handle diverse spectrum-to-structure tasks according to the available prior knowledge by bridging the retrieval and generation. It not only achieves state-of-the-art performance in analyzing theoretical Infrared and Raman spectra, but also outperform previous models at experimental data. Moreover, Vib2Mol demonstrates promising capabilities in predicting reaction products and sequencing peptides, enabling vibrational spectroscopy a real-time guide for autonomous scientific discovery workflows.111codes are available at https://github.com/X1nyuLu/vib2mol
@@ 1 Introduction
In this study, we propose a DL-based vibrational spectrum-to-molecular structure model (Vib2Mol) to flexibly address a variety of spectral annotation tasks according to the available prior knowledge. Vib2Mol adopts an encoder-decoder transformer architecture, and is trained with the strategy of multi-task learning and integrates a wide variety of spectrum-to-structure tasks into one versatile model. For a better evaluation, we compiled theoretical and experimental benchmarks, drawing upon both public datasets and our own Density functional theory (DFT) calculations. Overall, Vib2Mol not only achieves state-of-the-art performance on all benchmarks but also exhibits an overwhelming superiority when compared to mainstream methods. It further enhances the accuracy in interpreting spectra of reaction products and peptide sequencing as more knowledge of target molecules is introduced. This advancement demonstrates significant potential for in-situ intelligent analysis of dynamic chemical transformations and biological processes.
@@ 2.1 Multi-task learning framework: correlating vibrational spectrum and molecular structure
The workflow of Vib2Mol during pre-training, including alignment and generation phases, is illustrated in Figure 1. Vib2Mol adopts staged pre-training (SPT) as a fundamental training strategy. In the first stage, the alignment phase (Figure 1A) aims to bring the spectral and structural features of the same molecule as close as possible while separating the features of different molecules simultaneously. Spectra and molecular structures are represented as patch tokens and SMILES tokens, and then encoded into spectral and molecular embeddings by encoders, respectively. These two embeddings are effectively aligned through contrastive learning (CL), enabling cross-modal spectrum-structure retrieval. To further enhance retrieval performance, we deliberately selected hard negative samples—highly similar molecule-spectrum pairs—from each training batch. A matching loss was then utilized to guide the model in learning the subtle distinctions inherent in these challenging samples.
@@
Figure 1B depicts workflow of the second stage, including conditional generation and de novo generation of molecular structures. Conditional generation, i.e., predicting the masked molecular structure on the basis of the spectrum, draws on masked language modeling (MLM). Briefly, SMILES tokens, representing the molecular structure, are randomly masked by 45% and then processed by molecular encoders to generate molecular features. The spectral and molecular encoders were initialized by cloning the parameters from the alignment phase depicted in Figure 1A. These parameters were then subsequently fine-tuned specifically for the generation task to optimize performance. Then molecular decoders fused information from both masked molecular embeddings and spectral features, and predicted the to-be-determined tokens using cross-attention. Differently, de novo generation draws on language modeling (LM). SMILES tokens are sequentially masked from left to right and directly input into the molecular decoders sharing parameters with MLM. Guided by spectral features and previously generated SMILES sequences, the decoders can predict the next SMILES token from left to right until the entire sequence is complete. Note that the chemical formula is an optional input for Vib2Mol, as its provision unequivocally boosts performance. Nevertheless, Vib2Mol remains capable of achieving commendable results even in its absence, which will be discussed in Section 2.3.
@@
Figures 2A to 2E illustrate the workflow of Vib2Mol during application and inference. (1) For spectrum-spectrum retrieval (Figure 2A), instead of directly comparing spectral similarity by metrics such as Pearson correlation coefficient, the to-be-determined spectrum is encoded into an embedding vector, and the cosine similarity is calculated between this vector and the known spectral embedding vector in the database. (2) Spectrum-structure retrieval (Figure 2B) leverages cross-modal retrieval strategy. The spectrum-structure similarity is calculated between the embedding vector of the query spectrum and the molecular embedding vectors of known entities within the database. (3) For conditional generation (Figure 2C), Vib2Mol adopts the encoder-decoder architecture. Both the spectrum and partially masked molecular structure are encoded and then fused through the molecular decoder to generate the SMILES of the masked part. (4) For de novo generation (Figure 2D), Vib2Mol directly employs molecular decoders to sequentially predict each character of SMILES string on the basis of the encoded spectral features until a complete molecular structure is generated. When generating the next token, beam search (see Methods for details) is used to ensure the diversity of the results and then improve the generation performance. A re-ranking module (Figure 2E) is implemented to enhance both spectrum-structure retrieval and de novo generation. Candidate molecules were filtered based on the chemical formula, and then a pre-trained molecular encoder functions as a matching module, generating scores from a comprehensive evaluation of the query spectrum and candidate molecule features. Only candidates exhibiting high matching scores are subsequently selected as the final results. Even in the absence of chemical formula, the re-ranking module can still effectively sort candidates relying solely on its model-based scoring.
The model’s inherent parameter sharing and feature reusing allow it to address these four spectrum-to-structure tasks, without the need for additional fine-tuning or training. No wondering, spectrum-structure retrieval is more effective than spectrum-spectrum retrieval because it makes better use of molecular databases[15]. De novo generation offers more flexibility than conditional generation as it does not require a predefined molecular scaffold to predict side-chain structures. Therefore, for the spectrum-to-structure problem, spectrum-structure retrieval and de novo generation constitute the most versatile solutions, and will be the primary focus of following discussion.
@@
Figure 1: The framework of Vib2Mol for pretraining. (A) The alignment phase: spectra and molecular structures are represented as patch tokens and SMILES tokens, respectively. After processed by their encoders, spectral and molecular information are aligned by CL. Subsequently, hard negative samples are selected and employed to guide model in learning the subtle distinctions between these highly similar spectra or molecule samples. (B) The generation phase: for conditional generation, molecules are randomly masked 45% and encoded by the same molecular encoder used for spectrum-structure alignment. The molecular decoder fuses spectral information with molecular features and predicts masked tokens. For de novo generation, molecule is sequentially masked and directed input into the same molecular decoder as conditional generation without the prior encoding. Then, the decoder predicts the next token on the basis of previous information, spectral features and chemical formulae (if given).
Figure 2: The workflow of Vib2Mol for addressing different spectrum-to-structure tasks: (A) spectrum-spectrum retrieval, where only the spectral encoder is used to calculate the similarity between spectral pairs; (B) spectrum-structure retrieval, where spectra and molecules are encoded by their respective encoders to determine spectrum-structure similarity; (C) conditional generation, and (D) de novo generation, both following workflows during the stage of pretraining. (E) re-ranking module for refining retrieval and generation results. It initially filters candidates by chemical formula (if available), then uses a pre-trained molecular encoder to score them against the query spectrum. High-scoring candidates are finally selected as output.
@@ 2.2 State-of-the-art performance of Vib2Mol
To evaluate the performance of Vib2Mol on different spectrum-to-structure tasks and fairly compare it with current advanced models, five benchmarks are established. The benchmarks, which are described in detail in Methods, consist of three theoretical datasets (QM9S[16], VB-Mols, and VB-GEOM) and two experimental ones (SDBS[17] and NIST-IR[18]). Cumulatively, these benchmarks provide a total of 662,668 theoretical and 17,774 experimental spectra. There benchmarks enable a fair and comprehensive comparison with existing state-of-the-art methods, which can be broadly classified into three categories (see Methods for details): retrieval model (vibraCLIP[19]), generative models (PBSA[9], IR2Mol[20]), and comprehensive models capable of both retrieval and generation tasks (SMEN21, Vib2Mol). All benchmarking metrics are illustrated in Figure 2 and meticulously documented in Table S1-S5. Unless otherwise noted, Top-1 Recall (Recall@1) serves as the default metric throughout this section.
As depicted in Figure 3A, Vib2Mol demonstrated remarkable spectrum-structure retrieval performance on both the QM9S (98.11% for Raman, 96.63% for IR) and VB-Mols (94.66% for Raman, 93.38% for IR) benchmarks, outperforming vibraCLIP and performing on par with SMEN on QM9S (97.89% for Raman, 97.04% for IR) and VB-Mols (95.43% for Raman, 94.02% for IR). SMEN excels on theoretical benchmarks because it incorporates the precise molecular conformation as an extra input rather than simply using SMILES or 2D molecular graphs, which boosts its molecular representation and spectrum-retrieval performance. Despite this difference, Vib2Mol’s performance is still quite comparable.
The generalization of model was evaluated on the VB-GEOM benchmark, which differs from QM9S and VB-Mols in three ways. (1) As an out-of-distribution dataset, VB-GEOM provides a rigorous test of the robustness of models. (2) A different task, i.e., spectrum-spectrum retrieval is offered. (3) A zero-shot evaluation is performed, which means the model was pre-trained on the VB-Mols dataset and then directly evaluated on the VB-GEOM test set without any fine-tuning on VB-GEOM’s data. Vib2Mol demonstrated superior performance (77.54% for Raman, 75.33% for IR) among all models again, outperforming the second-place vibraCLIP (74.96% for Raman, 64.43% for IR) by a significant margin. In contrast, SMEN performed poorly (41.80% for Raman, 49.01% for IR), barely exceeding traditional cosine similarity and Pearson Correlation Coefficient, which represent the lower bound for this benchmark (Table S3).
When it comes to experimental benchmarks, all models were pre-trained on VB-Mols and then fine-tuned with the corresponding training set, and finally evaluated with the test set, given the inherent data scarcity of these experimental datasets. Vib2Mol demonstrated a substantially expanded lead in these benchmarks. Our model achieved the Recall@1 of 83.54%, 86.17%, and 90.43% on NIST-IR, SDBS-IR, and SDBS-Raman, respectively, which surpassing the metrics of vibraCLIP by almost 40 percentage points. Notably, SMEN was excluded since it could not be fine-tuned without molecular conformation data for target spectra in experimental benchmarks.
Figure 3B illustrates the de novo generation performance of various models across all benchmarks. Vib2Mol consistently achieved state-of-the-art performance. For the QM9S and VB-Mols benchmarks, Vib2Mol outperformed the second-best model by at least 10 percentage points. This performance gap was particularly evident in the VB-GEOM benchmark, which requires models to generate molecules from out-of-library spectra under zero-shot conditions. In this challenging scenario, Vib2Mol achieved a remarkable Recall@1 of 66.86% and 63.07% for Raman and infrared spectra, respectively. In contrast, the PBSA model yielded significantly lower metrics (50.90% and 47.32%), while the IR2Mol and SMEN fail to work on this scenario, resulting in a Recall@1 of 0%.The better performance of our model than others can be attributed to the meticulously designed pre-training strategy of Vib2Mol. Specifically, cross-modal alignment enables Vib2Mol to learn generalizable spectral and structural representations, while multi-task learning prevents overfitting to specific patterns in the training data.
We further examined the impact of integrating multi-modal spectral information on Vib2Mol performance. Figures 3C and 3D compare the retrieval and generation performance of Vib2Mol with multi-modal spectral input (Vib2Mol-MM) against Raman-only (Vib2Mol-Raman) and IR-only (Vib2Mol-IR) inputs. Details of processing multi-modal spectral signals can be found in supplementary information. A consistent trend emerged across all datasets: models based on Raman outperformed those based on IR, while models integrating both modalities significantly surpassed methods relying on single-modal input, which is consistent with related work[21]. Take de novo generation on the SDBS benchmark as an example, Vib2Mol-IR and Vib2Mol-Raman achieved Recall@1 scores of 52.84% and 56.03%, respectively, whereas, Vib2Mol-MM reached 68.44%, clearly demonstrating the benefit of multi-modal integration. The diverse molecular structural information provided by different spectra offers the model richer clues and a more integrated framework for constructing molecular structures or spectra. To further enhance the accuracy of structural elucidation, one could consider incorporating other spectral information from different modalities, such as NMR and MS, to provide more comprehensive perspectives on molecular structure features.
Figure 3: Performance evaluation of advanced deep learning models. (A) and (B) present a performance comparison of various models on spectrum-to-structure retrieval and de novo molecular generation, respectively. These evaluations were conducted on both theoretical (VB, QM9S) and experimental (NIST, SDBS) benchmarks. The impact of multi-modal spectral input on performance of Vib2Mol is further detailed in (C) for retrieval and (D) for generation.
2.3 Module optimization and synergistic integration in Vib2Mol
Vib2Mol consists of four modules: the retrieval (CL) module, matching (re-ranking) module, conditional generation (MLM) module, and de novo generation (LM) module. We explored the synergistic interaction among these modules across different tasks using a series of ablation experiments on the VB-Mols-Raman dataset.
For retrieval task, CL-only model achieves a Recall@1 of 88.46% at the beginning (Figure 4A). The subsequent integration of the matching loss alone elevated retrieval performance to 89.57%. This demonstrates that when joint losses are optimized, the matching module effectively serves as an auxiliary factor, significantly enhancing the primary retrieval task. The incorporation of chemical formulae for filtering inaccurately retrieved candidates further enhanced performance to 91.20%. Crucially, the final re-ranking step yielded the most substantial gain in Recall@1, achieving 94.66% (with chemical formula) and 93.20% (without). The consistent improvement underscores the efficacy of this comprehensive strategy.
For de novo generation (Figure 4B), data augmentation for SMILES strings emerged as the most significant strategy for improving metrics, elevating Recall@1 from 60.02% to 76.08%. This outcome is consistent with previous research20, largely because data augmentation prevents the model from overfitting to the syntactic patterns of standard SMILES, instead guiding it to learn the intrinsic correlation between spectral data and molecular structures. Following this, the introduction of SPT and MLM loss each contributed a modest improvement to the generation metrics. For more details on the ablation study related to MLM, please refer to the Supplementary Information. Crucially, the inclusion of chemical formulae then propelled the overall performance to a new level, rising from 77.49% to 81.93%. This substantial gain can be attributed to the chemical formula constraints, which effectively assist the model in determining elemental composition and overall unsaturation, thereby reducing uncertainty during generation. Furthermore, employing beam search enhanced the diversity of generated outputs, successfully preventing greedy decoding from converging on local optima. When a chemical formula was supplied, a chemical formula-based filter rigorously guaranteed the validity of the generated results through rule-based enforcement. Ultimately, the re-ranking module provided an additional boost to the overall Recall@1 of 86.74% (with chemical formula) and 82.59% (without).
Figure 4: Ablation studies of Vib2Mol on the VB-Mols-Raman dataset and visualization and statistical analysis of Vib2Mol representation learning. The performance contributions of different modules and hyperparameters are systematically assessed for three tasks: (A) retrieval and (B) de novo generation. (C) Alignment of spectral and structural embeddings, which illustrates the distribution of cosine similarities between spectrum and structure embeddings of the same molecule before and after training. (D) Performance on similar molecules categorized by functional groups, including retrieval and de novo generation tasks.
2.4 Spectrum-structure correlation captured by Vib2Mol
From the macro perspective, the effectiveness of Vib2Mol was explored on the test set of VB-Mols-Raman. Figure 4C visualizes the distribution of cosine similarities between spectrum-structure embeddings of individual molecule before and after training by Vib2Mol. The similarity distribution before training was concentrated around -0.023, and significantly increased to 0.86 after training, demonstrating the effective alignment between spectral and structural representations of each molecule upon training with Vib2Mol. To further investigate if our model truly understanding of the spectrum-structure relationship, we analyze its performance on failure cases. Figure S1 illustrates the Tanimoto similarity distributions between incorrectly generated or retrieved molecules and their targets. Among the incorrectly generated results, approximately 56% of the molecules had a Tanimoto similarity greater than 0.5 to their target molecule (0.5 is a common threshold for determining molecular similarity[14, 21]), with the overall distribution having an expectation of 0.53. Among the incorrectly retrieved molecules, about 47% had a Tanimoto similarity greater than 0.5, and the expectation for this distribution is 0.47. These results suggest that even when the outputs were incorrect, they still maintained a degree of structural similarity to their intended targets, such as misplacing a methyl group or an oxygen atom.
From the micro perspective, we explored the retrieval and generation performance of Vib2Mol on similar molecules by categorizing the test set into 15 common functional groups. Figure S2 presents the t-SNE visualization of the distribution of spectral and structural embeddings within these functional group subsets. In the majority of instances, the spectral and structural embeddings of the same molecule exhibited overlap, affirming their successful alignment. Following this, the model’s performance in spectrum-structure retrieval and de novo generation was assessed within each subclass (Figure 4D). For spectrum-structure retrieval, Vib2Mol achieved a weighted average Recall@1 of 90.75% across all functional groups. In de novo generation, Vib2Mol demonstrated a Recall@1 of 86.74% on the entire test set. Notably, when considering only the functional group accuracy of the generated molecules, the weighted average F1-score reached an impressive 97.87%. These results collectively underscore the robust capability of Vib2Mol in distinguishing among similar molecules.